# Advancing Time Series Wildfire Spread Prediction: Modeling Improvements and the WSTS+ Benchmark

### v2

@@ Abstract
Recent research has demonstrated the potential of deep neural networks (DNNs) to accurately predict wildfire spread on a given day based upon high-dimensional explanatory data from a single preceding day, or from a time series of $T$ preceding days. Here, we introduce a variety of modeling improvements that achieve state-of-the-art (SOTA) accuracy for both single-day and multi-day input scenarios, as evaluated on a large public benchmark for next-day wildfire spread, termed the WildfireSpreadTS (WSTS) benchmark. Consistent with prior work, we found that models using time-series input obtained the best overall accuracy. Furthermore, we create a new benchmark, WSTS+, by incorporating four additional years of historical wildfire data into the WSTS benchmark. Our benchmark doubles the number of unique years of historical data, expands its geographic scope, and, to our knowledge, represents the largest public benchmark for time-series-based wildfire spread prediction.
@@ 1 Introduction
Wildfires are a global cause of concern that have severe human, economical, and environmental impacts, with the average annual economic burden from wildfires falling between \$71.1 billion and \$347.8 billion [38]. Due to climate change, many regions are getting hotter and drier, which lengthens wildfire seasons and exacerbates their effect. In order to better manage, mitigate, and prevent wildfires, accurately predicting their spread is essential. In this work, we focus on the problem of next-day wildfire spread prediction, where we are provided with current and/or historical information about a particular wildfire, and then tasked with predicting its spatial extent on the following day.
Figure 1: The wildfire prediction models take as input a geospatial raster of several variables: vegetation, topography, and weather features, alongside the current day fire mask. We consider two prediction scenarios: one in which the model receives input features from only one preceding day, denoted $t-1$, and another in which we provide the model five previous days of features as input. In either case, the model must predict a binary mask indicating the extent of the fire on day $t$
A variety of approaches have been investigated to solve this problem, such as those based upon machine learning models [22, 4, 8], or physics-based and observationally-informed models [10, 1, 11]. In this work, however, we focus on a promising emerging class of techniques that utilize high-capacity machine learning models – namely deep neural networks (DNNs) – to predict wildfire spread using high-dimensional explanatory input data. These input data typically comprise a geospatial map of the current extent of the fire, as well as maps of explanatory features such as topography, climate, weather, and vegetation indices. Based upon these input data, the model is tasked with producing a geospatial map, or an image, reflecting the spatial extent of the fire on the following day, i.e. the spread of fire into regions that have not burned, and the extinction of fire in regions that have ceased burning. See Fig. 1 for an illustration.
DNN-based models have achieved impressive prediction accuracy, and have garnered substantial attention in recent years for wildfire modeling, including specifically next-day wildfire prediction. A variety of DNN-based models have been proposed to solve next-day prediction, including convolutional models [28, 5, 30], attention-based models such as transformers [37], and spatio-temporal models [31, 3]. Most existing research has focused on next-day prediction where only explanatory data from the current day is provided (day $t$ only, in Fig. 1). However, recent research found that models utilizing a time-series of $T$ days of historical data can achieve greater prediction accuracy [15], suggesting this as an important new direction in next-day wildfire prediction.
@@
In this work, we introduce a variety of modeling improvements to both the single-day ($T=1$) and time-series ($T>1$) input scenarios. Our results show that these improvements lead to substantial increases in accuracy over the existing state-of-the-art, as demonstrated on the WildfireSpreadTS (WSTS) benchmark [15], an extension of the well-studied Next Day Wildfire Spread benchmark [20] to support time-series prediction. We chose WSTS because it is the only publicly available benchmark for time-series wildfire prediction, and employs a rigorous and realistic 12-fold leave-one-year-out cross-validation. Our best model achieves a new SOTA on the WSTS benchmark by a large margin. Lastly, we introduce WSTS+, an extended benchmark for next-day wildfire spread, constructed by doubling the number of years of historical wildfire events in WSTS. The WSTS+ benchmark reveals an important emerging challenge in wildfire modeling: data heterogeneity. We summarize our contributions as follows:
  • SOTA Single-Day ($T=1$) Prediction. We introduce a variety of improvements to modeling single-day wildfire prediction, leading to $37\%$ improvement in performance over existing models.
  • SOTA Time-Series ($T>1$) Prediction. We introduce a variety of modeling improvements, resulting in a $28\%$ performance improvement in accuracy over existing models. We also find that our time-series models outperform single-day models, resulting in a new overall SOTA on wildfire prediction.
  • WSTS+: An expanded Public Benchmark for Next-Day Wildfire Modeling. Our proposed benchmark includes twice the years of historical data as its predecessor, WSTS, resulting, to our knowledge, in the largest public benchmark for time-series next-day wildfire spread prediction. It also reveals an emerging challenge for the wildfire modeling community: data heterogeneity.
The rest of the paper is structured as follows: we formulate our problem setting in Sec. 2, Sec. 3 reviews related works, Sec. 4 describes our adopted benchmark, we present some preliminaries in Sec. 5, Sec. 6 details our experimental setting and results, Sec. 7 introduces our expanded benchmark, and Sec. 8 summarizes our findings.
@@ 2 Problem Setting
In its general formulation, the goal of next-day wildfire spread prediction is to predict a wildfire’s spatial extent on some $t^{th}$ day, denoted $y(t)$, given explanatory data from one or more preceding days, denoted $x(t)$. We adopt the more specific settings of recent literature [15, 20], as illustrated in Fig. 1, which assume there are $T$ consecutive previous days of explanatory data, so that $x(t)=\{\tilde{x}(t-i)\}_{i=1}^{T}$, and each $\tilde{x}(t)$ comprises a geospatial raster, so that $\tilde{x}(t)\in\mathbb{R}^{H\times W\times C}$, where $H,W$ correspond to spatial dimensions, and $C$ represents the number of explanatory variables, which may include previous fire masks (e.g., $y(t)\subset\tilde{x}(t)$). The fire extent is encoded in a binary geospatial image, $y(t)\in\{0,1\}^{H\times W}$, where a value of one indicates the presence of a fire. Our goal is then to use a dataset of historical wildfire data to infer parameters, $\theta$ of a predictive model of the form $y(t)=f_{\theta}(x(t))$.
@@ 3 Related Works
To apply DNNs to fire spread prediction, recent research has framed the spread problem as a semantic segmentation one, and developed multiple datasets to support this framing. [20] created Next-Day Wildfire Spread, a dataset for mono-temporal fire spread prediction. They developed a custom convolutional autoencoder that takes as input various explanatory variables and outputs a binary mask indicating fire presence at each pixel. Similarly, [15] extended the dataset to include multi-temporal prediction, added more explanatory variables, and higher resolution fire masks. They achieved the best performance using a standard Unet model and a Unet model with temporal attention (UTAE) [14]. Other datasets aim to predict beyond next-day and forecast fire behavior several days in advance, e.g., [34] developed SeasFire Cube and trained Unet++ models [41] for medium-term fire prediction, between 8 and 64 days. [23] improved upon the collected data cube and found that the LSTM and ConvLSTM models outperformed the Fire Weather Index (FWI). In FireSight, [18] collected a dataset using remote sensing data from 20 datasets, and trained a 3D UNet model to model short-term fire hazard, between 3 and 8 days.
Other Approaches
Aside from segmentation, other formulations have been developed for modeling fire spread using deep learning (DL). [36] used a CNN-based Reinforcement Learning model that predicts the best action as burn or no burn given current conditions. [16] developed a probabilistic cellular automata model to simulate wildfire spread. In [26], the authors developed Sim2Real-Fire, a synthetic, high-resolution dataset for fire spread forecast and backtracking and outperformed the considered baselines using a custom Transformer model. We refer the readers to [40] for a more comprehensive review of DL for wildfire prediction.
@@
In contrast to most existing work (e.g., [24, 12, 37, 39]), we focus on utilizing a time-series of features for next-day wildfire spread prediction, which has been cited as an important emerging area of research [12, 15, 24]. Historically, time-series modeling has been challenging due to the lack of appropriate public datasets to train and evaluate models for this task. Recently, [15], building upon the work of [20], developed the first multi-temporal dataset for time-series prediction. Notably, they found that models using a time-series of input tend to outperform those using a single day, reinforcing the importance of this research direction.
@@ 4 The WSTS Benchmark
In this work, we employ the WildfireSpreadTS benchmark [15]. The dataset includes 607 wildfire events across the western United States between 2018 and 2021, totaling 13,607 daily multi-channel images. These 23 channels include data on active fires, weather, topography, and vegetation, resampled to a common resolution of 375 meters, providing a multi-modal and multi-temporal framework for modeling fire spread. A key feature of this benchmark is a rigorous 12-fold cross-validation evaluation procedure. Each fold of the cross-validation includes all wildfire events from a single year, so that the trained models are always evaluated on wildfire events from a previously unseen year, reflecting real-world use of wildfire prediction models.
5 Preliminaries
U-Net [35]
The U-Net is a widely-used architecture for segmentation, including for remote sensing applications [25]. There are now many variations of the U-Net, including modern variants that utilize attention (e.g., SwinUnet [6]). The U-Net typically includes three distinguishing architectural features: an encoder, $f_{\theta_{En}}$ with parameters $\theta_{En}$; a decoder, $f_{\theta_{De}}$ with parameters $\theta_{De}$; and skip connections between the encoder and decoder, as is illustrated in Fig. 2. The original U-Net in [35] employed an encoder with relatively few layers that were trained from scratch. It has been found, however, that utilizing larger pre-trained encoders (e.g., ResNet models [19]) can be beneficial, including specifically when applied to remote sensing data[21].
UTAE [14]
The UTAE was originally developed for satellite imagery, and is essentially a U-Net that has been modified to process a time-series of imagery. The UTAE encodes each entry in the time-series independently using a shared encoder, and then fuses the resulting embeddings from each day using a Lightweight Temporal Self-Attention (LTAE) block [13]. This process is illustrated in Fig. 2 in the context of our wildfire spread problem. Given a $T$-length time-series of input, the encoder produces a series of $T$ embeddings $z(t)=\{\tilde{z}(t-i)\}_{i=1}^{T}$ where $\tilde{z}(t)\in\mathbb{R}^{D_{4}\times\frac{H}{8}\times\frac{W}{8}}$ at the output of the last layer of the encoder, as illustrated in Fig. 2. Then the LTAE computes an attention mask, $a\in\mathbb{R}^{T\times\frac{H}{8}\times\frac{W}{8}}$, which is utilized to combine the $T$ embeddings. Before computing the temporal attention, LTAE adds a sinusoidal positional embedding, $p(\bar{t})$ to each input embedding, where $\bar{t}\in[1,365]$ is an integer representing the day of the year, and $p(\bar{t})$ maps $\bar{t}$ to a unique sinusoidal representation. This positional embedding is motivated by the original application of UTAE to agricultural segmentation, where the appropriate segmentation depends heavily upon the day of the year. Once the attention mask is computed, it is then upsampled, and applied to the encoder embeddings output at each resolution to collapse the temporal dimension. After all temporal dimensions are collapsed, a conventional U-Net-like decoder is applied to the collapsed embeddings.
Figure 2: Illustration of the U-Net and UTAE models, adapted from [14] to our wildfire problem. The U-Net consists of (a) an encoder and (b) a decoder, with skip connections connecting intermediate layers. For a U-Net we have $T=1$ and temporal fusion, part (c), would not be present. The UTAE is similar to the U-Net except $T>1$ and it includes Temporal Fusion, specifically with an LTAE.
@@ 6 Improving Wildfire Spread Prediction
In this section, we describe our methods for $T=1$ and $T=5$ scenarios, respectively, as well as experiments to support them (e.g., ablations). Results for our developed benchmark models are reported in Tab. 2, in terms of Average Precision (AP) using 12-fold leave-one-year-out cross-validation on the WSTS benchmark, following prior work [15]. Also following [15], we report model performance for three feature sets: vegetation features only (Veg), a combination of vegetation and topographic features (Multi), and all features (All), which includes additional weather forecast features. Full experimental details are provided in the supplemental information. Models in Tab. 2 with citations correspond to the three current best models on WSTS, as reported in [15]. All other models reported in Tab. 2 were developed in this work.
@@ 6.1 Single-Day Input ($T=1$)
The current $T=1$ SOTA utilizes a U-Net architecture with a ResNet-18 encoder, and is denoted Res18-Unet[15]. Therefore we focus our investigation on improving the Res18-Unet[15].
@@
We next describe our improvements to the Res18-Unet[15] at a high level. More details can be found in the supplement.
(i) Improved Encoders. We hypothesize that better performance may be obtained with larger encoders, or those that utilize attention mechanisms. Recent studies have indicated attention-based models to be superior to convolutional models for wildfire spread [24, 42, 37, 39], although these studies did not utilize the rigorous 12-fold leave-one-year out cross-validation in the WSTS benchmark that we use here. Therefore we investigate the use of a ResNet50 [19] encoder, as well as the attention-based SwinUnet-Tiny encoder [6].
(ii) Utilizing Pre-trained Parameters. Utilizing pre-trained weights to initialize training is a well-established technique to improve model accuracy in data-limited scenarios, such as the WSTS benchmark. We introduce pre-trained weights into each of the encoders that we consider (i.e., ResNet18, ResNet50, and SwinUnet), while the decoders are trained from scratch.
(iii) Improved Loss Functions. The existing SOTA Res18-Unet [15] is trained using weighted binary cross-entropy loss. However, it has been established that Jaccard/Dice losses are often superior alternatives for segmentation tasks [9], and focal loss has been shown effective for class imbalance [27] (the WSTS benchmark exhibits severe class imbalance), and for wildfire spread in particular [12]. Therefore we investigate and compare the aforementioned losses in our experiments.
(iv) Improved Hyperparameter Optimization. The existing SOTA Res18-Unet [15] was trained by selecting the model with the highest F1 score on the validation, however, all models on WSTS are evaluated utilizing the average precision (AP) metric [15]. We investigate aligning the validation and testing metrics by using AP for both.
For our experiments, we consider a U-Net with a ResNet-18 encoder (denoted Res18-Unet), and a ResNet-50 encoder (denoted Res50-Unet), and a SwinUnet-Tiny (denoted SwinUnet). For each of these models, we perform a grid search over all combinations of learning rates ($[1e-1,1e-2,1e-3,1e-4,1e-5]$), loss functions (BCE, Focal, Dice, Jaccard), and the use of pre-training or not (a binary choice). Following [15], we use a single fold of the 12-fold cross-validation, and only one of the three feature sets (the ”All” set) for this optimization. As discussed, in contrast to previous work, we utilize AP during validation to select the best models instead of F1. The focal loss has two hyperparameters: $\alpha$, set as the inverse frequency of positive class pixels, and $\gamma$, set to its default value of two.
@@
We found that pre-training was nearly always beneficial, and that Focal Loss usually yielded substantial improvements compared to our other candidate losses. Therefore, for the WSTS benchmark, we included both pre-training and focal loss in all three of our models: Res18-Unet, Res50-Unet, and SwinUnet. As an ablation study, Tab. 1 reports the performance of our Res18-Unet on the full WSTS benchmark, where we progressively remove each of our improvements to assess its impact. Our results indicate that each improvement is highly beneficial, or at least not significantly harmful.
@@
Tab. 2 reports the performance of our three models on the WSTS benchmark, compared to the best existing $T=1$ model, Res18-Unet[15]. Our Res18-Unet is identical to the Res18-Unet[15], except for our aforementioned modifications, and obtains substantially higher AP across all candidate input features considered: a 37% improvement on average.
@@
Our other two models, Res50-Unet and SwinUnet, also substantially outperform the existing Res18-Unet[15]. However, despite having approximately twice the number of trainable model parameters of Res18-Unet, we find that our model outperforms the two larger models in most cases. Our Res18-Unet also obtains the highest overall AP (0.468) for the $T=1$ models when utilizing the ”Multi” feature set, establishing a new SOTA on WSTS for $T=1$.
@@
Several recent studies have reported that large and/or attention-based models achieve SOTA accuracy for $T=1$ wildfire spread prediction [24, 42, 37, 39]. However, we find here that with simple improvements and appropriate optimization, Res18-Unet outperforms such models. We also suspect that the more rigorous (and potentially more real-world) leave-one-year-out cross-validation adopted by the WSTS benchmark may penalize more complex models for overfitting, compared to the Res18-Unet.
@@
In Fig. 3, we qualitatively evaluate our Res18-Unet and Res50-Unet against the Res18-Unet [15]. Each row corresponds to a fire event and the columns show the current fire, the next day label, and the predictions of each model. Yellow represents the fire extent, green shows correctly predicted burned areas, and red shows false positives. We observe that the original model tends to overpredict fire spread, leading to multiple red patches where no fire actually occurs. However, the model also underpredicts in areas where the fire spreads, capturing some, but not the full extent of the fire. On the other hand, we observe that our models make consistently more accurate predictions, with far fewer false positives, and slightly better matching green areas.
@@
Data Res18-Unet[15] 5 $0.333\pm 0.079$ $0.344\pm 0.076$ ${0.325\pm 0.108}$ 14.4M
@@
Feature UTAE[15] 5 $0.372\pm 0.088$ $0.350\pm 0.113$ $0.321\pm 0.135$ 1.1M
@@
Figure 3: Sample predictions made by the Res18-Unet [15], our Res18-Unet, and Res50-Unet. The two leftmost columns show the current fire spread $y(t-1)$ and the next-day label $y(t)$. True positive pixels are colored in green, while false positives are colored in red

### v3

# Improved Wildfire Spread Prediction with Time-Series Data and the WSTS+ Benchmark

@@ Abstract
Recent research has demonstrated the potential of deep neural networks (DNNs) to accurately predict wildfire spread on a given day based upon high-dimensional explanatory data from a single preceding day, or from a time series of $T$ preceding days. For the first time, we investigate a large number of existing data-driven wildfire modeling strategies under controlled conditions, revealing the best modeling strategies and resulting in models that achieve state-of-the-art (SOTA) accuracy for both single-day and multi-day input scenarios, as evaluated on a large public benchmark for next-day wildfire spread, termed the WildfireSpreadTS (WSTS) benchmark. Consistent with prior work, we found that models using time-series input obtained the best overall accuracy, suggesting this is an important future area of research. Furthermore, we create a new benchmark, WSTS+, by incorporating four additional years of historical wildfire data into the WSTS benchmark. Our benchmark doubles the number of unique years of historical data, expands its geographic scope, and, to our knowledge, represents the largest public benchmark for time-series-based wildfire spread prediction.
@@ 1 Introduction
Wildfires are a global cause of concern that have severe human, economical, and environmental impacts, with the average annual economic burden from wildfires falling between \$71.1 billion and \$347.8 billion [thomas2017costs]. In order to better manage, mitigate, and prevent wildfires, accurately predicting their spread is essential. In this work, we focus on the problem of next-day wildfire spread prediction, where we are provided with current and/or historical information about a particular wildfire, and then tasked with predicting its spatial extent on the following day.
Figure 1: The wildfire prediction models take as input a geospatial map of several variables: vegetation, topography, and weather features, alongside the current day fire mask. We consider two scenarios: one in which the model receives input features from one preceding day, denoted t1t-1, and one in which it receives input from five previous days
A variety of approaches have been investigated to solve this problem, such as those based upon machine learning models [khanmohammadi2022prediction, bot2022systematic, chetehouna2015predicting], or physics-based and observationally-informed models [Finney_1998, alexander2006evaluating, finney2006overview]. In this work, however, we focus on a promising emerging class of techniques that utilize high-capacity machine learning models – namely, deep neural networks (DNNs) – to predict wildfire spread using high-dimensional explanatory input data. These input data typically comprise a geospatial raster of the current extent of the fire, as well as explanatory features such as topography, climate, weather, and vegetation indices. Based upon these input data, the model is tasked with producing a geospatial map, or an image, reflecting the spatial extent of the fire on the following day. See Fig. 1 for an illustration.
A variety of DNN-based models have been proposed to solve next-day prediction, including convolutional models [liu2024fire, burge2023recurrent, marjani2024cnn], attention-based models such as transformers [shah2023wildfire], and spatio-temporal models [michail2024seasonal, bolt2022spatio]. One major limitation of most existing work is the lack of standardized evaluation wherein studies often utilize different datasets, model training and evaluation procedures, or compare to few other existing methods. Furthermore, most existing research has focused on next-day prediction where only explanatory data from the current day is input (e.g., only $\tilde{x}(t-1)$ in Fig. 1). However, recent research found that models utilizing a time-series of $T$ previous days of data can achieve greater prediction accuracy [gerard2023wildfirespreadts], suggesting this as an important new direction in next-day wildfire prediction.
@@
Our primary contributions are not a novel architecture, but rather, our first contribution is to perform a rigorous and controlled comparison of many existing approaches, both for single-day ($T=1$) and the time-series inputs ($T=5$). We also propose a novel temporal positional embedding. We compare all methods using the same public benchmark dataset, along with the same training, hyperparameter optimization, and evaluation procedures. We perform all experiments on the WildfireSpreadTS (WSTS) benchmark [gerard2023wildfirespreadts] because it is the only public benchmark for time-series wildfire prediction, it sufficiently large to support DNN training and evaluation, and in contrast to all other existing benchmarks, it employs a realistic 12-fold leave-one-year-out cross-validation. Our study reveals the best existing modeling strategies, resulting in substantial performance improvements over the current state-of-the-art (SOTA) for the WSTS benchmark.
Our second contribution is to introduce WSTS+, an extended benchmark for next-day wildfire spread, constructed by doubling the number of years of historical wildfire events in WSTS. By adopting WSTS+, we can double the size of our training datasets. We conduct experiments that reveal additional historical training data, either in WSTS or WSTS+, yields little improvement in the accuracy of wildfire models. We discover and analyze significant cross-year domain shift, a critical challenge for the field.
The rest of the paper is structured as follows: we formulate our problem setting in Sec. 2, Sec. 3 reviews related works, Sec. 4 describes WSTS, Sec. 5 details our experiments, Sec. 6 introduces WSTS+, and Sec. 7 concludes.
@@ 2 Problem Setting
In its general formulation, the goal of next-day wildfire spread prediction is to predict a wildfire’s spatial extent on some $t^{th}$ day, denoted $y(t)$, given explanatory data from one or more preceding days, denoted $x(t)$. We adopt the more specific settings of recent literature [gerard2023wildfirespreadts, huot2022next], as illustrated in Fig. 1, which assume there are $T$ consecutive previous days of explanatory data, so that $x(t)=\{\tilde{x}(t-i)\}_{i=1}^{T}$, and each $\tilde{x}(t)$ comprises a geospatial raster, so that $\tilde{x}(t)\in\mathbb{R}^{H\times W\times C}$, where $H,W$ correspond to spatial dimensions, and $C$ represents the number of explanatory variables, which may include previous fire masks (e.g., $y(t)\subset\tilde{x}(t)$). The fire extent is encoded in a binary geospatial image, $y(t)\in\{0,1\}^{H\times W}$, where a value of one indicates the presence of a fire. Our goal is then to use a dataset of historical wildfire data to infer parameters, $\theta$, of a predictive model of the form $y(t)=f_{\theta}(x(t))$.
@@ 3 Related Works
DNN-based segmentation has received growing attention due to its accuracy, enabled by the recent development of large datasets of historical fire data. [huot2022next] created Next-Day Wildfire Spread, a large and public dataset for next-day spread prediction, and used it to train a custom deep segmentation model. Concurrently, [prapas2022deep] developed SeasFire Cube and trained Unet++ models [zhou2018unet++] for medium-term fire prediction, between 8 and 64 days. [kondylatos2022wildfire] improved upon the collected data cube and found that the LSTM and ConvLSTM models outperformed the Fire Weather Index (FWI). In FireSight, [gottfriedsenfiresight] collected a dataset using remote sensing data from 20 datasets, and trained a 3D UNet model to model short-term fire hazard, between 3 and 8 days. Overall, most work has been done using U-Net architectures and their variants, and many authors [li2024wildfire, fitzgerald2023paying, shah2023wildfire, xiao2024wildfire] have recently reported that attention-based U-Nets achieve greater accuracy. We investigate the SwinUnet[liu2021swin] in our study as a widely used and therefore representative example of such models.
@@
In contrast to the existing work discussed above, we also focus on time-series input for spread prediction, which has been cited as an important emerging direction [fitzgerald2023paying, gerard2023wildfirespreadts, li2024wildfire]. Historically, time-series modeling has been challenging due to the lack of appropriate public datasets to train and evaluate models for this task. Recently, [gerard2023wildfirespreadts] extended the Next-Day Wildfire Spread dataset from [huot2022next] to be suitable for time-series prediction, and achieved their best overall next-day predictions using a time-series model, termed UTAE [garnot2021panoptic].
Other DL Approaches
Aside from a segmentation formulation, researchers have also investigated, for example, reinforcement learning [ross2021being], probabilistic cellular automata [ghosh2024fire], and synthetic data approaches [lisim2real]. We refer readers to [xu2024wildfire] for a review of DL for wildfire prediction.
@@ 4 The WSTS Benchmark
In this work, we employ the WildfireSpreadTS benchmark [gerard2023wildfirespreadts]. The dataset includes 607 wildfire events across the western United States between 2018 and 2021, totaling 13,607 daily multi-channel images. These 23 channels include data on active fires, weather, topography, and vegetation, resampled to a common resolution of 375 meters, providing a multi-modal and multi-temporal framework for modeling fire spread. A key feature of this benchmark is a rigorous 12-fold cross-validation evaluation procedure. Each fold of the cross-validation includes all wildfire events from a single year, so that the trained models are always evaluated on wildfire events from a previously unseen year, reflecting real-world use of wildfire prediction models.
@@ 5 Improving Wildfire Spread Prediction
In this section, we describe our methods for $T=1$ and $T=5$ scenarios, respectively, as well as experiments to support them (e.g., ablations). Results for our developed benchmark models are reported in Tab. 2, in terms of Average Precision (AP) using 12-fold leave-one-year-out cross-validation on the WSTS benchmark, following prior work [gerard2023wildfirespreadts]. Also following [gerard2023wildfirespreadts], we report model performance for three feature sets: vegetation features only (Veg), a combination of vegetation and topographic features (Multi), and all features (All), which includes additional weather forecast features. Detailed descriptions of our models, alongside full experimental details, are provided in Sec. 8 in the supplement. Key deployment information (FLOPs, GPU memory requirements, training/inference times) can be found in Sec. 9.1. Models in Tab. 2 with citations correspond to the three current best models on WSTS, as reported in [gerard2023wildfirespreadts]. All other models reported in Tab. 2 were developed in this work.
@@ 5.1 Single-Day Input ($T=1$)
The current $T=1$ SOTA utilizes a U-Net architecture with a ResNet-18 encoder, and is denoted Res18-Unet[gerard2023wildfirespreadts]. Therefore, we focus our investigation on improving the Res18-Unet[gerard2023wildfirespreadts].
@@
We next describe the investigated improvements to the Res18-Unet[gerard2023wildfirespreadts] at a high level. More details can be found in Sec. 8 of the supplement.
(i) Encoders. Better performance may be obtained with larger encoders or those with attention mechanisms. Recent studies have indicated that attention-based models may be superior to convolutional models for wildfire spread [li2024wildfire, zou2023attention, shah2023wildfire, xiao2024wildfire]. We investigate a ResNet50 [he2016deep] encoder, as well as the attention-based SwinUnet-Tiny encoder [cao2022swin] and SegFormer-B2 [xie2021segformer].
(ii) Utilizing Pre-trained Parameters. Utilizing pre-trained weights to initialize training is a well-established technique to improve model accuracy, including in remote sensing [iglovikov2018ternausnet]. We investigate pre-trained weights for each of the encoders that we consider (i.e., ResNet18, ResNet50, SwinUnet, and SegFormer), while the decoders are trained from scratch.
(iii) Improved Loss Functions. The existing SOTA Res18-Unet [gerard2023wildfirespreadts] is trained using weighted binary cross-entropy loss. However, it has been established that Jaccard/Dice losses are often superior alternatives for segmentation tasks [eelbode2020optimization], and focal loss has been shown effective for class imbalance [lin2017focal] (the WSTS benchmark exhibits severe class imbalance), and for wildfire spread in particular [fitzgerald2023paying]. Therefore we investigate and compare the aforementioned losses in our experiments.
(iv) Improved Hyperparameter Optimization. The existing SOTA Res18-Unet [gerard2023wildfirespreadts] was trained by selecting the model with the highest F1 score on the validation; however, all models on WSTS are evaluated utilizing the average precision (AP) metric [gerard2023wildfirespreadts]. We investigate aligning the validation and testing metrics by using AP for both.
For our experiments, we consider a U-Net with a ResNet-18 encoder (denoted Res18-Unet), a ResNet-50 encoder (denoted Res50-Unet), a SwinUnet-Tiny (denoted SwinUnet), and a SegFormer-B2 (denoted SegFormer). For each of these models, we perform a grid search over all combinations of learning rates ([1e1,1e2,1e3,1e4,1e5][1e-1,1e-2,1e-3,1e-4,1e-5]), loss functions (BCE, Focal, Dice, Jaccard), and the use of pre-training or not (a binary choice). Following [gerard2023wildfirespreadts], we use a single fold of the 12-fold cross-validation, and only one of the three feature sets (the ”All” set) for this optimization. As discussed, in contrast to previous work, we utilize AP during validation to select the best models instead of F1. The focal loss has two hyperparameters: $\alpha$, set as the inverse frequency of positive class pixels, and $\gamma$, set to its default value of two.
@@
We found that pre-training was nearly always beneficial, and that Focal Loss usually yielded substantial improvements compared to our other candidate losses. Therefore, for the WSTS benchmark, we included both pre-training and focal loss in all our models: Res18-Unet, Res50-Unet, SwinUnet, and SegFormer. As an ablation study, Tab. 1 reports the performance of our Res18-Unet on the full WSTS benchmark, where we progressively remove each of our improvements to assess its impact. Our results indicate that each improvement is highly beneficial, or at least not significantly harmful.
@@
Tab. 2 reports the performance of our models on the WSTS benchmark, compared to the best existing $T=1$ model, Res18-Unet[gerard2023wildfirespreadts]. Our Res18-Unet is identical to the Res18-Unet[gerard2023wildfirespreadts], except for our aforementioned modifications, and obtains substantially higher AP across all input features considered: a 37% improvement on average. We find that these improvements are statistically significant using a Wilcoxon signed-rank test, and report the results in Sec. 9.3.
@@
Our other models, Res50-Unet, SwinUnet, and SegFormer also substantially outperform the existing Res18-Unet[gerard2023wildfirespreadts]. However, despite having approximately twice the number of trainable model parameters of Res18-Unet, we find that our model outperforms the three larger models in most cases. Our Res18-Unet also obtains the highest overall AP (0.468) for the $T=1$ models when utilizing the ”Multi” feature set, establishing a new SOTA on WSTS for $T=1$.
@@
Several recent studies have reported that large and/or attention-based models achieve SOTA accuracy for $T=1$ wildfire spread prediction [li2024wildfire, zou2023attention, shah2023wildfire, xiao2024wildfire]. However, we find here that with simple improvements and appropriate optimization, Res18-Unet outperforms such models. In Sec. 9.2 and Tab. 8, we test the hypothesis that the more rigorous (and potentially more real-world) leave-one-year-out cross-validation adopted by the WSTS benchmark may penalize more complex models for overfitting. Yet, we find that using a random cross-validation still allows the Res18-Unet to outperform attention-based models, suggesting its superiority even under different evaluation scenarios.
@@
In Fig. 2, we qualitatively evaluate our Res18-Unet and Res50-Unet against the Res18-Unet [gerard2023wildfirespreadts]. Each row corresponds to a fire event, and the columns show the current fire, the next-day label, and the predictions of each model. Yellow represents the fire extent, green shows correctly predicted burned areas, and red shows false positives.
We observe that the original model tends to overpredict fire spread, leading to multiple red patches where no fire actually occurs. However, the model also underpredicts in areas where the fire spreads, capturing some, but not the full extent of the fire. On the other hand, we observe that our models make consistently more accurate predictions, with far fewer false positives, and slightly better matching green areas.
We provide additional examples in Fig. 11, through Fig. 18 of the best and worst predictions made by the Res18-Unet for each testing year, and we analyze these results in Sec. 9.4 and Sec. 9.5. Overall, we find that the model struggles with very small, newly ignited, or displaced fires, while achieving high accuracy on larger, more consolidated fires, with performance positively correlated with fire size across years.
@@
SegFormer 1 $0.433\pm 0.080$ $0.436\pm 0.083$ $0.423\pm 0.087$ 27.5M
Data Res18-Unet[gerard2023wildfirespreadts] 5 $0.333\pm 0.079$ $0.344\pm 0.076$ $0.325\pm 0.108$} 14.4M
@@
SegFormer 5 $0.439\pm 0.081$ $0.436\pm 0.085$ $0.430\pm 0.082$ 27.7M
Feature UTAE[gerard2023wildfirespreadts] 5 $0.372\pm 0.088$ $0.350\pm 0.113$ $0.321\pm 0.135$ 1.1M
@@
Figure 2: Sample predictions made by the Res18-Unet [gerard2023wildfirespreadts], our Res18-Unet, and Res50-Unet. The two leftmost columns show the current fire spread y(t1)y(t-1) and the next-day label $y(t)$. True positive pixels are colored in green, while false positives are colored in red